Given f (x)= x/x-3 - 8/x-2 ; solve f (x) = 0
Answer by Cromlix(4381)

Hi there,
f (x)= x/x-3 - 8/x-2
Common denominator
= (x - 3)(x - 2)
Cross multiply
x(x - 2) -8(x - 3)
x^2 - 2x -8x + 24
x^2 - 10x + 24
x^2 - 10x + 24 = 0
(x - 6)(x - 4) = 0
x - 6 = 0
x = 6
x - 4 = 0
x = 4
x = 6,4.
for f(x) = 0
Hope this helps :-)